# Jian-hao Kanga, Zhun Goub and Nan-jing Huangc

### v1

@@ 6 Conclusions
$+0.5\left(b^{2}\hat{q}^{2}(s)+\hat{\pi}^{2}(s)(X^{\hat{u}}(s))^{2}\sigma^{2}\right)H_{xx}(s,X^{\hat{u}}(s),M_{1}^{\hat{u}}(s))$
$+\left(X^{\hat{u}}(s)-\alpha M_{1}^{\hat{u}}(s)-e^{-\alpha h}M_{2}^{\hat{u}}(s)\right)H_{m_{1}}(s,X^{\hat{u}}(s),M_{1}^{\hat{u}}(s))\bigg)\mathrm{d}s\bigg].$ (C.49)
We emphasize that equations (C) and (C) represent general results valid for all $U,H\in C^{1,2,1}(\mathcal{D})$ satisfying the condition $(A1)$, independent of both equations (3.9) and (3.10). The terminal condition $U(T,x,m_{1})=H(T,x,m_{1})$ together with equations (3.8), (3.11) and (C.46)-(C) collectively yield
$U(t,x,m_{1})$ $=H(t,x,m_{1})=\int(\varphi^{\gamma})^{-1}\left(Y^{\gamma}(t,x,m_{1})\right)\mathrm{d}\Gamma(\gamma)$
$=\int(\varphi^{\gamma})^{-1}\left(y^{\hat{u},\gamma}(t,x,m_{1},m_{2})\right)\mathrm{d}\Gamma(\gamma)=J^{\hat{u}}(t,x,m_{1},m_{2}).$ (C.50)
@@
$J^{u_{\varepsilon}}(t,x,m_{1},m_{2})=$ $\mathbb{E}_{t}\left[J^{u_{\varepsilon}}\left(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon)\right)\right]$
$-\mathbb{E}_{t}\left[\int(\varphi^{\gamma})^{-1}\left(y^{u_{\varepsilon},\gamma}(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon))\right)\mathrm{d}\Gamma(\gamma)\right]$
$+\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{u_{\varepsilon},\gamma}(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon))\right]\right)\mathrm{d}\Gamma(\gamma).$
@@
$J^{u_{\varepsilon}}\left(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon)\right)=U\left(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon)\right).$
@@
$y^{u_{\varepsilon},\gamma}\left(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon)\right)=y^{\hat{u},\gamma}\left(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon)\right).$
@@
$J^{u_{\varepsilon}}(t,x,m_{1},m_{2})=$ $\mathbb{E}_{t}\left[U(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]$
$-\mathbb{E}_{t}\left[\int(\varphi^{\gamma})^{-1}\left(y^{\hat{u},\gamma}(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon))\right)\mathrm{d}\Gamma(\gamma)\right]$
$+\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{\hat{u},\gamma}(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon))\right]\right)\mathrm{d}\Gamma(\gamma).$ (C.51)
@@
$\mathcal{A}^{u}U(t,x,m_{1})-\mathcal{A}^{u}H(t,x,m_{1})+\int\iota^{\gamma}\left(y^{\hat{u},\gamma}(t,x,m_{1},m_{2})\right)\mathcal{A}^{u}y^{\hat{u},\gamma}(t,x,m_{1},m_{2})\mathrm{d}\Gamma(\gamma)\leq 0.$ (C.52)
Given the continuity of the admissible reinsurance and investment strategy $u(t,x,m_{1},m_{2})$, we first apply Lemma 3.2 for arbitrary $\varepsilon>0$ and then take the limit $\varepsilon\rightarrow 0$ to establish the following identities
$\mathbb{E}_{t}\left[U\left(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon)\right)\right]-U(t,x,m_{1})$
$=$ $\varepsilon\left\{\left[(A+\pi(t,x,m_{1},m_{2})(\mu-r))x+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q(t,x,m_{1},m_{2})\right]U_{x}(t,x,m_{1})\right$.
$+U_{t}(t,x,m_{1})+0.5\left(b^{2}q^{2}(t,x,m_{1},m_{2})+\pi^{2}(t,x,m_{1},m_{2})x^{2}\sigma^{2}\right)U_{xx}(t,x,m_{1})$
$\left.+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)U_{m_{1}}(t,x,m_{1})\right\}+o(\varepsilon)$ (C.53)
@@
$\mathbb{E}_{t}\left[H\left(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon)\right)\right]-H(t,x,m_{1})$
$=$ $\varepsilon\left\{\left[(A+\pi(t,x,m_{1},m_{2})(\mu-r))x+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q(t,x,m_{1},m_{2})\right]H_{x}(t,x,m_{1})\right$.
$+H_{t}(t,x,m_{1})+0.5\left(b^{2}q^{2}(t,x,m_{1},m_{2})+\pi^{2}(t,x,m_{1},m_{2})x^{2}\sigma^{2}\right)H_{xx}(t,x,m_{1})$
$\left.+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)H_{m_{1}}(t,x,m_{1})\right\}+o(\varepsilon)$ (C.54)
@@
$\mathbb{E}_{t}\left[y^{\hat{u},\gamma}\left(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon)\right)\right]-y^{\hat{u},\gamma}(t,x,m_{1},m_{2})$
$=$ $\varepsilon\left\{\left[(A+\pi(t,x,m_{1},m_{2})(\mu-r))x+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q(t,x,m_{1},m_{2})\right]y^{\hat{u},\gamma}_{x}(t,x,m_{1},m_{2})\right$.
$+y^{\hat{u},\gamma}_{t}(t,x,m_{1},m_{2})+0.5\left(b^{2}q^{2}(t,x,m_{1},m_{2})+\pi^{2}(t,x,m_{1},m_{2})x^{2}\sigma^{2}\right)y^{\hat{u},\gamma}_{xx}(t,x,m_{1},m_{2})$
$\left.+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)y^{\hat{u},\gamma}_{m_{1}}(t,x,m_{1},m_{2})\right\}+o(\varepsilon),$ (C.55)
where $o(\varepsilon)$ represents a higher-order infinitesimal and in the equation (C), we have used the fact derived from the equation (C.46) that $y^{\hat{u},\gamma}(t,x,m_{1},m_{2})$ is independent of $m_{2}$. Then, from the equation (C), one has
$\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{\hat{u},\gamma}\left(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon)\right)\right]\right)\mathrm{d}\Gamma(\gamma)$
$\quad-\int(\varphi^{\gamma})^{-1}\left(y^{\hat{u},\gamma}(t,x,m_{1},m_{2})\right)\mathrm{d}\Gamma(\gamma)$
$=$ $\varepsilon\left\{\int\iota^{\gamma}\left(y^{\hat{u},\gamma}(t,x,m_{1},m_{2})\right)\bigg(y^{\hat{u},\gamma}_{t}(t,x,m_{1},m_{2})+\big[(A+\pi(t,x,m_{1},m_{2})(\mu-r))x\right$.
$+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q(t,x,m_{1},m_{2})\big]y^{\hat{u},\gamma}_{x}(t,x,m_{1},m_{2})$
$+0.5\left(b^{2}q^{2}(t,x,m_{1},m_{2})+\pi^{2}(t,x,m_{1},m_{2})x^{2}\sigma^{2}\right)y^{\hat{u},\gamma}_{xx}(t,x,m_{1},m_{2})$
$+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)y^{\hat{u},\gamma}_{m_{1}}(t,x,m_{1},m_{2})\bigg)\mathrm{d}\Gamma(\gamma)\bigg\}+o(\varepsilon).$ (C.56)
@@
$\mathbb{E}_{t}\left[U(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]-U(t,x,m_{1})-\mathbb{E}_{t}\left[H(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]$
$\quad+H(t,x,m_{1})+\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{\hat{u},\gamma}(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon))\right]\right)\mathrm{d}\Gamma(\gamma)$
$\quad-\int(\varphi^{\gamma})^{-1}\left(y^{\hat{u},\gamma}(t,x,m_{1},m_{1})\right)\mathrm{d}\Gamma(\gamma)\leq o(\varepsilon).$
@@
$U(t,x,m_{1})\geq$ $\mathbb{E}_{t}\left[U(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]-\mathbb{E}_{t}\left[H(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]$
$+\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{\hat{u},\gamma}(t+\varepsilon,X^{u}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon))\right]\right)\mathrm{d}\Gamma(\gamma)+o(\varepsilon)$
$=$ $J^{u_{\varepsilon}}(t,x,m_{1},m_{2})+o(\varepsilon)$
and so $J^{\hat{u}}(t,x,m_{1},m_{2})\geq J^{u_{\varepsilon}}(t,x,m_{1},m_{2})+o(\varepsilon)$. Thus, $\underset{\varepsilon\rightarrow 0}{\liminf}\frac{J^{\hat{u}}(t,x,m_{1},m_{2})-J^{u_{\varepsilon}}(t,x,m_{1},m_{2})}{\varepsilon}\geq 0,$ which establishes $\hat{u}$ as an equilibrium reinsurance and investment strategy. Therefore, the identity $V(t,x,m_{1})=U(t,x,m_{1})$ is established, thereby completing the proof. ∎
Appendix D D Proof of Proposition 4.1
@@
Since A=rBCA=r-B-C, $C=\beta e^{-\alpha h}$ and $Be^{-\alpha h}=(\alpha+A+\beta)C$, we can rewrite $\hat{q}(t)$ as
$\hat{q}(t)=\frac{a\eta_{2}}{b^{2}\mathbb{E}[\gamma]}\exp\left\{\left[-r+\frac{\beta}{1+\beta}\left(r+\alpha+e^{-\alpha h}-1\right)\right](T-t)\right\}.$
Therefore, we have

### v2

@@ 7 Conclusions
$+0.5\left(b^{2}\hat{q}^{2}(s)+\hat{\pi}^{2}(s)(X^{\hat{u}}(s))^{2}\sigma^{2}S^{2\delta}_{1}(s)\right)H_{xx}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))$
$+\mu S_{1}(s)H_{s_{1}}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))+0.5\sigma^{2}S^{2\delta+2}_{1}(s)H_{s_{1}s_{1}}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))$
$+\left(X^{\hat{u}}(s)-\alpha M_{1}^{\hat{u}}(s)-e^{-\alpha h}M_{2}^{\hat{u}}(s)\right)H_{m_{1}}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))$
$+\hat{\pi}(s)X^{\hat{u}}(s)\sigma^{2}S^{2\delta+1}_{1}(s)H_{xs_{1}}(s,X^{\hat{u}}(s),S_{1}(s),M_{1}^{\hat{u}}(s))\bigg)\mathrm{d}s\bigg].$ (C.59)
We emphasize that equations (C) and (C) represent general results valid for all $U,H\in C^{1,2,2,1}(\mathcal{D})$ satisfying the condition $(A1)$, independent of both equations (3.9) and (3.10). The terminal condition $U(T,x,s_{1},m_{1})=H(T,x,s_{1},m_{1})$ together with equations (3.8), (3.11) and (C.56)-(C) collectively yield
$U(t,x,s_{1},m_{1})$ $=H(t,x,s_{1},m_{1})=\int(\varphi^{\gamma})^{-1}\left(Y^{\gamma}(t,x,s_{1},m_{1})\right)\mathrm{d}\Gamma(\gamma)$
$=\int(\varphi^{\gamma})^{-1}\left(y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{2})\right)\mathrm{d}\Gamma(\gamma)=J^{\hat{u}}(t,x,s_{1},m_{1},m_{2}).$ (C.60)
@@
$J^{u_{\varepsilon}}(t,x,s_{1},m_{1},m_{2})=$ $\mathbb{E}_{t}\left[J^{u_{\varepsilon}}\left(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon)\right)\right]$
$-\mathbb{E}_{t}\left[\int(\varphi^{\gamma})^{-1}\left(y^{u_{\varepsilon},\gamma}(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon))\right)\mathrm{d}\Gamma(\gamma)\right]$
$+\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{u_{\varepsilon},\gamma}(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon))\right]\right)\mathrm{d}\Gamma(\gamma).$
@@
$J^{u_{\varepsilon}}\left(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon)\right)=U\left(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon)\right).$
@@
$y^{u_{\varepsilon},\gamma}\left(t+\varepsilon,X^{u_{\varepsilon}}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u_{\varepsilon}}(t+\varepsilon),M_{2}^{u_{\varepsilon}}(t+\varepsilon)\right)=y^{\hat{u},\gamma}\left(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon)\right).$
@@
$J^{u_{\varepsilon}}(t,x,s_{1},m_{1},m_{2})=$ $\mathbb{E}_{t}\left[U(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]$
$-\mathbb{E}_{t}\left[\int(\varphi^{\gamma})^{-1}\left(y^{\hat{u},\gamma}(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon))\right)\mathrm{d}\Gamma(\gamma)\right]$
$+\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{\hat{u},\gamma}(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon))\right]\right)\mathrm{d}\Gamma(\gamma).$ (C.61)
@@
$\mathcal{A}^{u}U(t,x,s_{1},m_{1})-\mathcal{A}^{u}H(t,x,s_{1},m_{1})+\int\iota^{\gamma}\left(y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{2})\right)\mathcal{A}^{u}y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{2})\mathrm{d}\Gamma(\gamma)\leq 0.$ (C.62)
Given the continuity of the admissible reinsurance and investment strategy $u(t,x,s_{1},m_{1},m_{2})$, we first apply Lemma 3.2 for arbitrary $\varepsilon>0$ and then take the limit $\varepsilon\rightarrow 0$ to establish the following identities
$\mathbb{E}_{t}\left[U\left(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon)\right)\right]-U(t,x,s_{1},m_{1})$
$=$ $\varepsilon\left\{\left[(A+\pi(t,x,s_{1},m_{1},m_{2})(\mu-r))x+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q(t,x,s_{1},m_{1},m_{2})\right]U_{x}(t,x,s_{1},m_{1})\right$.
$+U_{t}(t,x,s_{1},m_{1})+0.5\left(b^{2}q^{2}(t,x,s_{1},m_{1},m_{2})+\pi^{2}(t,x,s_{1},m_{1},m_{2})x^{2}\sigma^{2}s^{2\delta}_{1}\right)U_{xx}(t,x,s_{1},m_{1})$
$+\mu s_{1}U_{s_{1}}(t,x,s_{1},m_{1})+0.5\sigma^{2}s^{2\delta+2}_{1}U_{s_{1}s_{1}}(t,x,s_{1},m_{1})+\pi(t,x,s_{1},m_{1},m_{2})x\sigma^{2}s^{2\delta+1}_{1}$
$\left.\times U_{xs_{1}}(t,x,s_{1},m_{1})+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)U_{m_{1}}(t,x,s_{1},m_{1})\right\}+o(\varepsilon)$ (C.63)
@@
$\mathbb{E}_{t}\left[H\left(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon)\right)\right]-H(t,x,s_{1},m_{1})$
$=$ $\varepsilon\left\{\left[(A+\pi(t,x,s_{1},m_{1},m_{2})(\mu-r))x+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q(t,x,s_{1},m_{1},m_{2})\right]H_{x}(t,x,s_{1},m_{1})\right$.
$+H_{t}(t,x,s_{1},m_{1})+0.5\left(b^{2}q^{2}(t,x,s_{1},m_{1},m_{2})+\pi^{2}(t,x,s_{1},m_{1},m_{2})x^{2}\sigma^{2}s^{2\delta}_{1}\right)H_{xx}(t,x,s_{1},m_{1})$
$+\mu s_{1}H_{s_{1}}(t,x,s_{1},m_{1})+0.5\sigma^{2}s^{2\delta+2}_{1}H_{s_{1}s_{1}}(t,x,s_{1},m_{1})+\pi(t,x,s_{1},m_{1},m_{2})x\sigma^{2}s^{2\delta+1}_{1}$
$\left.\times H_{xs_{1}}(t,x,s_{1},m_{1})+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)H_{m_{1}}(t,x,s_{1},m_{1})\right\}+o(\varepsilon)$ (C.64)
@@
$\mathbb{E}_{t}\left[y^{\hat{u},\gamma}\left(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon)\right)\right]-y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{2})$
$=$ $\varepsilon\left\{\left[(A+\pi(t,x,s_{1},m_{1},m_{2})(\mu-r))x+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q(t,x,s_{1},m_{1},m_{2})\right]y^{\hat{u},\gamma}_{x}(t,x,s_{1},m_{1},m_{2})\right$.
$+y^{\hat{u},\gamma}_{t}(t,x,s_{1},m_{1},m_{2})+0.5\left(b^{2}q^{2}(t,x,s_{1},m_{1},m_{2})+\pi^{2}(t,x,s_{1},m_{1},m_{2})x^{2}\sigma^{2}s^{2\delta}_{1}\right)y^{\hat{u},\gamma}_{xx}(t,x,s_{1},m_{1},m_{2})$
$+\mu s_{1}y^{\hat{u},\gamma}_{s_{1}}(t,x,s_{1},m_{1},m_{2})+0.5\sigma^{2}s^{2\delta+2}_{1}y^{\hat{u},\gamma}_{s_{1}s_{1}}(t,x,s_{1},m_{1},m_{2})+\pi(t,x,s_{1},m_{1},m_{2})x\sigma^{2}s^{2\delta+1}_{1}$
$\left.\times y^{\hat{u},\gamma}_{xs_{1}}(t,x,s_{1},m_{1},m_{2})+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)y^{\hat{u},\gamma}_{m_{1}}(t,x,s_{1},m_{1},m_{2})\right\}+o(\varepsilon),$ (C.65)
where $o(\varepsilon)$ represents a higher-order infinitesimal and we use the fact derived from the equation (C.56) that $y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{2})$ is independent of $m_{2}$. Then, from the equation (C), one has
$\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{\hat{u},\gamma}\left(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon)\right)\right]\right)\mathrm{d}\Gamma(\gamma)$
$\quad-\int(\varphi^{\gamma})^{-1}\left(y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{2})\right)\mathrm{d}\Gamma(\gamma)$
$=$ $\varepsilon\left\{\int\iota^{\gamma}\left(y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{2})\right)\bigg(y^{\hat{u},\gamma}_{t}(t,x,s_{1},m_{1},m_{2})+\big[(A+\pi(t,x,s_{1},m_{1},m_{2})(\mu-r))x\right$.
$+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q(t,x,s_{1},m_{1},m_{2})\big]y^{\hat{u},\gamma}_{x}(t,x,s_{1},m_{1},m_{2})$
$+0.5\left(b^{2}q^{2}(t,x,s_{1},m_{1},m_{2})+\pi^{2}(t,x,s_{1},m_{1},m_{2})x^{2}\sigma^{2}s^{2\delta}_{1}\right)y^{\hat{u},\gamma}_{xx}(t,x,s_{1},m_{1},m_{2})$
$+\mu s_{1}y^{\hat{u},\gamma}_{s_{1}}(t,x,s_{1},m_{1},m_{2})+0.5\sigma^{2}s^{2\delta+2}_{1}y^{\hat{u},\gamma}_{s_{1}s_{1}}(t,x,s_{1},m_{1},m_{2})+\pi(t,x,s_{1},m_{1},m_{2})x\sigma^{2}s^{2\delta+1}_{1}$
$\times y^{\hat{u},\gamma}_{xs_{1}}(t,x,s_{1},m_{1},m_{2})+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)y^{\hat{u},\gamma}_{m_{1}}(t,x,s_{1},m_{1},m_{2})\bigg)\mathrm{d}\Gamma(\gamma)\bigg\}+o(\varepsilon).$ (C.66)
@@
$\mathbb{E}_{t}\left[U(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]-U(t,x,s_{1},m_{1})-\mathbb{E}_{t}\left[H(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]$
$\quad+H(t,x,s_{1},m_{1})+\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{\hat{u},\gamma}(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon))\right]\right)\mathrm{d}\Gamma(\gamma)$
$\quad-\int(\varphi^{\gamma})^{-1}\left(y^{\hat{u},\gamma}(t,x,s_{1},m_{1},m_{1})\right)\mathrm{d}\Gamma(\gamma)\leq o(\varepsilon).$
@@
$U(t,x,s_{1},m_{1})\geq$ $\mathbb{E}_{t}\left[U(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]-\mathbb{E}_{t}\left[H(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon))\right]$
$+\int(\varphi^{\gamma})^{-1}\left(\mathbb{E}_{t}\left[y^{\hat{u},\gamma}(t+\varepsilon,X^{u}(t+\varepsilon),S_{1}(t+\varepsilon),M_{1}^{u}(t+\varepsilon),M_{2}^{u}(t+\varepsilon))\right]\right)\mathrm{d}\Gamma(\gamma)+o(\varepsilon)$
$=$ $J^{u_{\varepsilon}}(t,x,s_{1},m_{1},m_{2})+o(\varepsilon)$
and so $J^{\hat{u}}(t,x,s_{1},m_{1},m_{2})\geq J^{u_{\varepsilon}}(t,x,s_{1},m_{1},m_{2})+o(\varepsilon)$. Thus, $\underset{\varepsilon\rightarrow 0}{\liminf}\frac{J^{\hat{u}}(t,x,s_{1},m_{1},m_{2})-J^{u_{\varepsilon}}(t,x,s_{1},m_{1},m_{2})}{\varepsilon}\geq 0,$ which establishes $\hat{u}$ as an equilibrium reinsurance and investment strategy. Therefore, the identity $V(t,x,s_{1},m_{1},m_{2})=U(t,x,s_{1},m_{1})$ is established, thereby completing the proof. ∎
Appendix D D Proof of Theorem 4.1
@@
We aim to confirm that the candidate equilibrium reinsurance and investment strategy (4.22) is admissible and that the assumptions of Theorem 3.1 are met.
Setting $\widetilde{Z}(t)=\frac{1}{4\delta^{2}}S^{-2\delta}_{1}(t)$, we derive that $\mathrm{d}\widetilde{Z}(t)=\widetilde{\kappa}(\widetilde{\theta}-\widetilde{Z}(t))\mathrm{d}t+\widetilde{\sigma}\sqrt{\widetilde{Z}(t)}\mathrm{d}W_{2}(t)$, where $\widetilde{\kappa}=2\mu\delta$, $\widetilde{\theta}=\frac{2\delta+1}{8\mu\delta^{2}}\sigma^{2}$ and $\widetilde{\sigma}=-\sigma$. It is easy to see that $2\widetilde{\kappa}\widetilde{\theta}>\widetilde{\sigma}^{2}$, which implies that $\widetilde{Z}(t)$ is almost surely non-negative. Moreover, by the proof process of Lemma 1 in [27], one has $\mathbb{E}_{t}\big[\sup\limits_{s\in[t,T]}|\widetilde{Z}(s)|^{p}\big]<\infty$ for $p\in[1,\infty)$.
Now, when the expression (4.22) is substituted into the SDDE (2.3), it leads to